LOCAL RESTAURANT GUIDE

The Dragon’s Lair Café Hill University Center – Breakfast and lunch and dinner/Chick Fil-A/Deli Bar Fast Food Restaurants near the Campus Numerous fast food franchises are located adjacent to the campus on 6th Avenue South between 15th and 18th Street. Other Restaurants near the Campus

• Anthony’s – 2131 7th Avenue South – Italian – 324-1215

• Burly Earl’s – 2109 7th Avenue South – Sandwiches, Billards – 322-5848

• La Cocina – 2111 7th Avenue South – Mexican – 252-7626 – Mon-Thu/11am-9 pm – Fri/11am-10pm • Dreamland – 1427 14th Avenue South – BarBQ Ribs – 933-2133 – Mon-Sun

• Fishmarket Restaurant – 611 21st Street South – Seafood – 322-3330

• Lucy’s Coffee and Tea – 2007 University Blvd – 328-2007 – Mon-Fri/7am-5pm

• New York Bagel Café & Deli – 801 20th Street South – open @7:30am-4pm

• The Pita Shop Café – 1106 12th Street South – Lebanese – 328-2749 –

• Mon-Fri/11am-10pm

• Purple Onion – 1717 10th Avenue South – Pita Sandwiches, Gyro – 933-2424 – Mon-Fri/9am-2am • Ted’s – 328 12th Street South – breakfast/Lunch – 324-2911 – Mon-Fri – Breakfast/6-10am; Lunch

10am-2:30pm

• Tracy’s – 729 20th Street South – Cafeteria style – Mon-Fri – Breakfast/7-10am; Lunch 10:45am-

2:15pm Restaurants and Delis in Five Points South

• Back Alley – 1312 Cobb Lane – Dinner/Sandwiches/Music – 933-6211 – Tue-Sat/11:am-11:pm

• Cobb Lane Restaurant – 1 Cobb Lane – Reservations – 933-0462 – Mon-Sat/11am-2:30am

• Cosmo’s Pizza – 2012 Magnolia Avenue – 930-9971 – Mon-Sat/11:am-10:pm

• Golden Temple – 1901 11th Avenue South – Vegetarian – Lunch only – 933-8933 – Mon-Sat/11am-2:30pm

• Highland’s – 2011 11th Avenue South – Reservations – 939-1400 – Tue-Sat/6 pm-10 pm

• Mancha’s – 1207 20th Street South – Mexican bar/restaurant – 939-3304 – Mon-Thurs/11:am-11:pm • The Mill – 1035 20th Street South – Bakery, Brewery and Restaurant – Mon-Fri/11am-6:30pm

• The Original Pancake House – 1931 11th Avenue South – 933-8837 – Mon-Fri/6am-2:am • Ranelli’s – 1225 20th Avenue South – Italian deli – 933-6983 – Mon-Fri/11am-6:pm

• Surin West – 1918 11th Avenue South – Thai food, Lunchtime – 324-1928 – Mon-Fri/11am-5:30pm • Zydeco – 2001 15th Avenue South – Fast Food/Bar/Live music – 933-1032 – Mon-Fri/4pm-Until

Outlying restaurants worth a trip • Bogue’s – 3028 Clairmont Avenue – breakfast, lunch – 254-9780 – Mon-Fri/6:am-10:30pm

• Bottega Café – 2242 Highland Avenue South – Casual – 933-2001 – Mon-Fri/10:am-10:30pm

• Bottega Restaurant – 2240 Highland Avenue South – Fancy – Reservations – 933-1000

• Taj India – 2226 Highland Avenue South – Great Indian food, Lunch buffet – 939-3805

• Hot n Hot - Fish Club – 2180 11th Court South – Fancy – 933-5474

• Salvatore’s – 720 Old Town Road, Vestavia Hills – Italian Food – 822-7310

• Stix Restaurant Japanese Steakhouse – 3250 Galleria Circle, Hoover – 982-3070